Baichung’s father is 26 years younger than Baichung’s grandfather and 29 years older than Baichung. The sum of the ages of all the three is 135 years. What is the age of each one of them?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the age of Baichung’s father be x.
Then, the age of Baichung’s grandfather = (x+26)
and, Age of Baichung = (x-29) According to the question, x + (x+26) + (x-29) = 135
⇒3x+26-29=135
⇒3x-3=135
⇒3x=135+3
⇒3x=138
⇒x=138/3
⇒x=46
Age of Baichung’s father = x = 46
Age of Baichung’s grandfather = (x+26) = 46 + 26 = 72
Age of Baichung=(x-29)=46-29=17

Thanks